OpenStudy (anonymous):

Word Problem. a. A college student has set aside $240 for the rest of the school year to use the coin-operated laundry facility in his dormitory. Each time he uses the machines, it costs $7.50. Choose the equation that represents the amount remaining in his fund,f, after he has done laundry x times. Find the amount remaining in the fund after12 trips to the laundry facility.

OpenStudy (anonymous):

@ texaschix101......

OpenStudy (texaschic101):

f = 240 - 7.5x f = 240 - 7.5(12) f = 240 - 90 f = 150 <-- amount left after using laundry mat 12 times

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

anytime :)